OpenStudy (anonymous):

If i have an 18 inch pizza, what is it's interior angle and arc length of the slices? What is the area of each slice?

jimthompson5910 (jim_thompson5910):

How many slices are there?

OpenStudy (anonymous):

i need one for 12 slices and for 14 slices :( please explain?

jimthompson5910 (jim_thompson5910):

The diameter is 18 so the circumference is C = pi*d C = pi*18 C = 3.14*18 C = 56.52 So the circumference is roughly 56.52 inches

jimthompson5910 (jim_thompson5910):

you divide this circumference by 12 to get the arc length of each slice

OpenStudy (anonymous):

so i do the same thing for 14? divide this same circumference?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

oh ok, Thank you so much!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

but wait i need to find the interior angle.. how do i do that with the circumference?

OpenStudy (anonymous):

and how do i find the area of each slice? do i use the triangle geometry formula?

jimthompson5910 (jim_thompson5910):

Area of One Slice = (1/n)*pi*r^2 where n = number of slices and r = radius

jimthompson5910 (jim_thompson5910):

So for 12 slices, n = 12 in this case, r = 9, so Area of One Slice = (1/n)*pi*r^2 Area of One Slice = (1/12)*3.14*9^2 Area of One Slice = 21.195 square inches (roughly)

OpenStudy (anonymous):

so whats the formula for interior angles? or do you not know, cause i dont.

jimthompson5910 (jim_thompson5910):

i think you mean central angle central angle = 360/n where n = number of slices

OpenStudy (anonymous):

the question asked for the interior angle, and i dont remember the formulas v.v

jimthompson5910 (jim_thompson5910):

well if you have a regular polygon, then the interior angle is 180(n-2)/n but this is a circle

OpenStudy (anonymous):

i guess i would do that with however many slices for sides maybe....?

jimthompson5910 (jim_thompson5910):

yeah i guess so, it seems like an odd choice of wording if you ask me

OpenStudy (anonymous):

^agreed

OpenStudy (anonymous):

thank you anyway you've been a huge help!

jimthompson5910 (jim_thompson5910):

yw